I have a test tomorrow and i'm having trouble understanding this question. Please help me.
The perimeter of a triangle is 17 centimeters. One side is 3 centimeters longer than the shortest side, whereas the third side is 2 centimeters shorter than twice the shortest side. Find the lengths of the three sides

Answer by addingup(3677)

A triangle has three sides, they are a, b, and c (lower case, upper case is used for angles)
Now just turn the words into algebra, like this:
One side: a
shortest side: b
third side: c
and it says that:
a = b+3
c = 2b-2
so the problem references all sides to b. Piece of cake!
-----------------------------------------------------
a+b+c = 17 substitute like this:
b+3+b+2b-2 = 17
4b+1 = 17
4b = 16
b = 4 so now we know that b = 4, lets find the others:
a = b+3 = 4+3 = 7
c = 2b-2 = 2(4)-2 = 8-2 = 6
Now that we have all our values, let's check and make sure we've got the right numbers:
a+b+c = 17
7+4+6 = 17 add the left side and you'll see I've got the correct answer.
:
Happy learning,
John